# Multi: Multimodal Understanding Leaderboard with Text and Images

### v3

@@ 4.10 What is the main cause of the error?
To explore the primary causes of errors, we analyzed 165 questions (one-third of Multi-Elite), using predictions generated by the 8 models evaluated in § 4.9. To get more detailed insights, we focused on outputs generated with the CoT prompt, which helps reveal patterns and challenges across different error types. The detailed distribution of all errors is shown in Figure 9, and we give five example cases with error analysis for several selected models as shown in F. We also asked human experts to identify key abilities required to solve each question, such as comprehension of text and images, mathematical calculation, logical reasoning, knowledge recall, and others. The distribution of these abilities is shown in Figure 8, allowing us to compare the alignment between the skills demanded by the questions and the following observed error patterns.
We summarize these errors into five primary categories, with the following observations:
Alignment Errors (11.1%): A significant proportion of errors cause by models failing to follow CoT (Qwen2-VL-7B in Figure 15 and Figure 18) or formatting instructions (InternVL2-8B in Figure 18), highlighting difficulties in instruction compliance. Some models also rejected questions outright, potentially due to divergences in human alignment objectives. Notably, Qwen2-VL-7B and 72B exhibited higher rejection rates when prompted with CoT but proceeded to provide direct answers, as discussed in § 4.9.
@@
Image Understanding Errors (24.4%): Much different from text understanding, errors in image summarization and detailed analysis account for 15% to 30% of total errors. Since 107 (65%) out of the 165 questions analyzed involve images, it demonstrates the challenge of handling image inputs. Similar to text errors, detail-level image comprehension caused more errors, as it is the bottleneck for all models to handle the AAT questions just like the one in Figure 15. There are cases in which models often struggle to locate and focus on specific image elements as mentioned in the text (GPT-4o in Figure 18), sometimes resulting in visual hallucinations. Enhanced strategies for analyzing and synthesizing image details are crucial to address this issue.
@@ 4.11 Takeaways
  • Models generally perform better on questions requiring shorter answers, i.e. SA $>$ MA $>$ FB $>$ OP. The results of MA Acc. emphasize the importance of balancing recall and precision, where less capable MLLMs failed to provide all correct choices. (Table 6 in § 4.4)
@@
  • The inclusion of Multi-Extend is likely to help improve performance on models with longer window sizes. However, it might also have adverse effects due to multiple reasons. (Figure 6 in § 4.6)
@@
  • Errors in model performance are primarily driven by failures in knowledge recall, logical reasoning, and image understanding. Additionally, alignment and instruction-following issues underscore challenges in maintaining adherence to complex prompts, particularly in chain-of-thought tasks. The observed error types match closely with the skill requirements identified by human annotators. (Figure 8 and Figure 9 in § 4.10)
@@ Appendix A Limitations and Future Work
  • Metrics for evaluating blank-filling, open-ended writing and others: Our evaluation primarily uses exact match, which might be overly stringent for assessing MLLMs’ true capabilities. Assessing open-ended writing tasks that require complex knowledge and reasoning is still a challenge. We also have nearly 100 questions that do not belong to traditional categories, such as questions requiring geographic drawing, and the evaluation of them will be even more challenging. Now that only a few studies Wang et al. (2023) involve human evaluation, developing automatic and reliable methods remains an open research area.
@@
Appendix B Models and Experiment Cost
@@
FDU MOSS Sun et al. (2023) 2023-04-19 16B Weight T moss-moon-003-sft
AISpeech DFM-2.0 Chen et al. (2022) 2024-01-16 70B Weight T dfm-2.0-70b-preview
Alibaba Qwen2-7B qwe (2024) 2024-06-04 7B Weight T Qwen2-7B-Instruct
PJLab InternLM2.5-7B Cai et al. (2024) 2024-06-27 7B Weight T internlm2_5-7b-chat
[…]
OpenBMB VisCPM Hu et al. (2023) 2023-06-28 10B Weight SI VisCPM-Chat
LinkSoul-AI Chinese-LLaVA LinkSoul-AI (2023) 2023-06-30 7B Weight One Chinese-LLaVA-Cllama2
Alibaba Qwen-VL Bai et al. (2023a) 2023-09-25 7B Weight MI Qwen-VL-Chat
01-ai Yi-VL-34B 01.ai (2023) 2023-11-22 34B Weight One Yi-VL-34B
OpenGVLab InternVL-1.1 Chen et al. (2024) 2024-01-24 19B Weight One InternVL-Chat-V1-1
OpenGVLab InternVL2-8B Zhe et al. (2024) 2024-07-03 8B Weight MI InternVL2-8B
OpenBMB MiniCPM-V-2.6 Yao et al. (2024) 2024-08-05 8B Weight MI MiniCPM-V-2_6
Alibaba Qwen2-VL-7B Wang et al. (2024) 2024-08-28 7B Weight MI Qwen2-VL-7B-Instruct
Alibaba Qwen2-VL-72B Wang et al. (2024) 2024-08-28 72B Weight MI Qwen2-VL-72B-Instruct
OpenAI ChatGPT OpenAI (2022) 2023-11-06 - API T gpt-3.5-turbo-1106
OpenAI GPT-4 OpenAI (2023a) 2023-11-06 - API T gpt-4-1106-preview
Google Gemini Pro Team (2023) 2023-12-07 - API T gemini-1.0-pro-001
[…]
Google Gemini Vision Pro Team (2023) 2023-12-07 - API MI gemini-1.0-pro-vision
OpenAI GPT-4o-mini OpenAI (2024a) 2024-07-18 - API MI gpt-4o-mini-2024-07-18
OpenAI GPT-4o (0806) OpenAI (2024b) 2024-08-06 - API MI gpt-4o-2024-08-06
Google Gemini 1.5 Flash Team (2023) 2024-08-27 - API MI gemini-1.5-flash-exp-0827
Anthropic Claude 3.5 Sonnet Anthropic $-2$ API MI claude-3-5-sonnet-20241022
OpenAI GPT-4o (1120) OpenAI (2024b) 2024-11-20 - API MI gpt-4o-2024-11-20
Google Gemini 1.5 Pro Team (2023) 2024-11-21 - API MI gemini-exp-1121
@@
A standard evaluation using GPT-4o with image inputs (without CoT reasoning) requires approximately 4.5 million prompt tokens and 150,000 completion tokens, amounting to a total cost of around \$25. Running on a single A100(80G) GPU, models with 7–8B parameters complete the benchmark in 1–2 hours, while a 70B+ model takes approximately 8 hours on 2-4 A100 GPUs. CoT experiments demand significantly more resources due to the larger token output. To facilitate more practical and faster usage, we recommend researchers prioritize testing on MULTI-Elite, which requires only around \$1 and a few minutes to complete.
Appendix C Prompts
@@
Appendix D Data Selection Algorithm
@@
$L_{q}=\left(a\times\begin{bmatrix}\mathcal{H}(L_{q,\#\text{characters in question}})\\ \mathcal{H}(L_{q,\#\text{characters in answer}})\\ \mathcal{H}(L_{q,\#\text{characters in analysis}})\end{bmatrix}\right.+\left.b\times\begin{bmatrix}\mathcal{H}(L_{q,\#\text{images in question}})\\ \mathcal{H}(L_{q,\#\text{images in answer}})\\ \mathcal{H}(L_{q,\#\text{images in analysis}})\end{bmatrix}\right)^{\top}\begin{bmatrix}1.0\\ 0.1\\ 0.5\end{bmatrix}$
@@
In the formula above, we use a harmonic mean function $\mathcal{H}$ to normalize content length $L_{q,i}$ of each target value $i$ within each knowledge piece $k$ [^5] Note that for those questions without knowledge information, we simply use an empty string as a keyword.
@@
$Pr[\text{pick up }q]=\begin{cases}1&\text{ , for }q\text{ s.t. }L_{q,k}[0]\\ p&\text{ , if }q=1\text{ , for }q\text{ of }L_{q,k}[1:m]\text{ or }L_{q,k}[-m:]\\ p\frac{N_{k}-2m}{\#\text{questions of }k}&\text{ , otherwise}\end{cases}$
Appendix E Data Process and Annotation
@@
E.1 Data Pre-process
The raw data range from HTML, photocopy, hand script, and plain text, and we conduct pre-processing to reduce the load of further annotation. We remove most HTML tags indicating irrelevant content of the question such as alignment, color, etc. We reserve tags for underlines (<u> </u>), and we transfer several tagged styles including bold, italic, and tabular data into markdown format. For some complex tables that cannot be well converted, we save them as a screenshot picture after rendering with HTML.
For photocopy and hand script, we adopt OCR tools to convert text content, crop images, and figures, and integrate them into markdown. We further transcript most of the math functions and chemistry structures into  format, with a small portion remaining as images.
E.2 Data Annotation
Figure 11: A screenshot for the main page of the data annotation platform.
We develop an online platform for data annotation stage. The platform consists of text boxes for editing contents and regions for rendering the text to see the final appearance of the data as shown in Figure 11. We employ skilled human annotators and involve them as authors, primarily undergraduate students from top universities in China familiar with exam quizzes and markdown rules, to undertake this comprehensive task covering various aspects from formatting to semantic analysis:
@@ E.2 Data Annotation
[…]
In Figure 12, we show several examples of complex formation and modification during the data annotation stage. The markdown, , and HTML format code are remained for better format clarity.
@@
E.3 Data Post-process
To collect more challenging data for our benchmark, we adopt several data post-process strategies:
@@ E.3 Data Post-process
  • Distillation. This is completed during our annotation process. We reduce assistance information so that the answer must depend on more detailed analysis (refer to example (4) in Figure 13). In this way, we greatly enhance question difficulty.
  • Transformation. We randomly modify the questions such as from single-choice to blank-filling (refer to example (2) in Figure 13), or convert certain kinds of single-choice questions into multiple-choice ones (refer to example (1) and (5) in Figure 13). Lots of single-choice questions have a list of options and the choices are presented as the combination of those options where only one is correct. We transform those questions into multiple-choice questions where the options become new choices and the correct answer corresponds to the combinations. In this way we successfully increase the scale of multiple-choice questions, improving the diversity of the questions.
@@
Appendix F Case study
@@
Figure 15: Error analysis of a SI-SA question.
Figure 16: Error analysis of a SI-SA question.
Figure 17: Error analysis of a MI-SA question.
Figure 18: Error analysis of a SI-OP question.
Figure 19: Error analysis of a NI-MA question.

### v4

@@ 4.10 What is the main cause of the error?
We asked human experts to identify key abilities required to solve each question, such as comprehension of text and images, mathematical calculation, logical reasoning, knowledge recall, and others. The distribution of these abilities is shown in Table 9, allowing us to compare the alignment between the skills demanded by the questions and the following observed error patterns.
To explore the primary causes of errors, we analyzed 165 questions (one-third of Multi-Elite), using predictions generated by the 8 models evaluated in § 4.9. To get more detailed insights, we focused on outputs generated with the CoT prompt, which helps reveal patterns and challenges across different error types. The detailed distribution of all errors is shown in Figure 9, and we give five example cases with error analysis for several selected models as shown in G.
We summarize these errors into six primary categories, with the following observations:
Alignment Errors (11.1%): A significant proportion of errors caused by models failing to follow CoT (Qwen2-VL-7B in Figure 15 and Figure 18) or formatting instructions (InternVL2-8B in Figure 18), highlighting difficulties in instruction compliance. Some models also rejected questions outright, potentially due to divergences in human alignment objectives. Notably, Qwen2-VL-7B and 72B exhibited higher rejection rates when prompted with CoT but proceeded to provide direct answers, as discussed in § 4.9.
@@
Image Understanding Errors (24.4%): Much different from text understanding, errors in image summarization and detailed analysis account for 15% to 30% of total errors. Since 107 (65%) out of the 165 questions analyzed involve images, it demonstrates the challenge of handling image inputs. Similar to text errors, detail-level image comprehension caused more errors, as it is the bottleneck for all models to handle the AAT questions, just like the one in Figure 15. There are cases in which models often struggle to locate and focus on specific image elements as mentioned in the text (GPT-4o in Figure 18), sometimes resulting in visual hallucinations. Enhanced strategies for analyzing and synthesizing image details are crucial to address this issue.
@@ 4.11 Takeaways
  • Models generally perform better on questions requiring shorter answers, i.e., SA >> MA >> FB >> OP. The results of MA Acc. emphasize the importance of balancing recall and precision, where less capable MLLMs failed to provide all correct choices. (Table 6 in § 4.4)
@@
  • The inclusion of Multi-Extend is likely to help improve performance on models with longer window sizes. However, it might also have adverse effects due to multiple reasons. (Table 7 in § 4.6)
@@
  • Errors in model performance are primarily driven by failures in knowledge recall, logical reasoning, and image understanding. Additionally, alignment and instruction-following issues underscore challenges in maintaining adherence to complex prompts, particularly in chain-of-thought tasks. The observed error types match closely with the skill requirements identified by human annotators. (Table 9 and Figure 9 in § 4.10)
@@ 5 Conclusion
[…]
This work was funded by National Natural Science Foundation of China Projects (Grant Nos. 92370206, U23B2057, and 62120106006), National Science and Technology Major Project (Grant No. 2023ZD0120703), and Shanghai Municipal Science and Technology Projects (Grant Nos. 2021SHZDZX0102 and 25X010202846). The authors would like to thank Kunyao LAN and Lei PAN for their contribution during the revision process.
@@ Appendix A Limitations and Future Work
  • Metrics for evaluating blank-filling, open-ended writing, and others: Our evaluation primarily uses exact match, which might be overly stringent for assessing MLLMs’ true capabilities. Assessing open-ended writing tasks that require complex knowledge and reasoning is still a challenge. We also have nearly 100 questions that do not belong to traditional categories, such as questions requiring geographic drawing, and the evaluation of them will be even more challenging. Now that only a few studies [59] involve human evaluation, developing automatic and reliable methods remains an open research area.
@@
Appendix B Several Useful Links
Evaluation Code of Multi: https://github.com/OpenDFM/MULTI-Benchmark
Appendix C Models and Experiment Cost
@@
FDU MOSS [1] 2023-04-19 16 B Weight T moss-moon-003-sft
AISpeech DFM-2.0 [2] 2024-01-16 70 B Weight T dfm-2.0-70b-preview
Alibaba Qwen2-7B [3] 2024-06-04 7 B Weight T Qwen2-7B-Instruct
PJLab InternLM2.5-7B [4] 2024-06-27 7 B Weight T internlm2_5-7b-chat
THU VisualGLM [15] 2023-05-17 6 B Weight SI visualglm-6b
OpenBMB VisCPM [16] 2023-06-28 10 B Weight SI VisCPM-Chat
LinkSoul-AI Chinese-LLaVA [17] 2023-06-30 7 B Weight One Chinese-LLaVA-Cllama2
Alibaba Qwen-VL [18] 2023-09-25 7 B Weight MI Qwen-VL-Chat
01-ai Yi-VL-34B [19] 2023-11-22 34 B Weight One Yi-VL-34B
OpenGVLab InternVL-1.1 [20] 2024-01-24 19 B Weight One InternVL-Chat-V1-1
OpenGVLab InternVL2-8B [21] 2024-07-03 8 B Weight MI InternVL2-8B
OpenBMB MiniCPM-V-2.6 [22] 2024-08-05 8 B Weight MI MiniCPM-V-2_6
Alibaba Qwen2-VL-7B [23] 2024-08-28 7 B Weight MI Qwen2-VL-7B-Instruct
Alibaba Qwen2-VL-72B [23] 2024-08-28 72 B Weight MI Qwen2-VL-72B-Instruct
OpenAI ChatGPT [5] 2023-11-06 - API T gpt-3.5-turbo-1106
OpenAI GPT-4 [6] 2023-11-06 - API T gpt-4-1106-preview
Google Gemini Pro [7] 2023-12-07 - API T gemini-1.0-pro-001
OpenAI GPT-4V [24] 2023-11-06 - API MI gpt-4-1106-vision-preview
Google Gemini Vision Pro [7] 2023-12-07 - API MI gemini-1.0-pro-vision
OpenAI GPT-4o-mini [25] 2024-07-18 - API MI gpt-4o-mini-2024-07-18
OpenAI GPT-4o (0806) [26] 2024-08-06 - API MI gpt-4o-2024-08-06
Google Gemini 1.5 Flash [7] 2024-08-27 - API MI gemini-1.5-flash-exp-0827
Anthropic Claude 3.5 Sonnet [27] 2024-10-22 - API MI claude-3-5-sonnet-20241022
OpenAI GPT-4o (1120) [26] 2024-11-20 - API MI gpt-4o-2024-11-20
Google Gemini 1.5 Pro [7] 2024-11-21 - API MI gemini-exp-1121
@@
A standard evaluation using GPT-4o with image inputs (without CoT reasoning) requires approximately 4.5 million prompt tokens and 150k completion tokens, amounting to a total cost of around \$25. Running on a single A100 (80 G) GPU, models with 7–8 B parameters complete the benchmark in 1–2 h, while a 70 B+ model takes approximately 8 h on 2-4 A100 GPUs. CoT experiments demand significantly more resources due to the larger token output. To facilitate more practical and faster usage, we recommend researchers prioritize testing on Multi-Elite, which requires only around \$1 and a few minutes to complete.
Appendix D Prompts
@@
Appendix E Data Selection Algorithm
@@
[…]
@@
In the formula above, we use a harmonic mean function $\mathcal{H}$ to normalize content length $L_{q,i}$ of each target value $i$ within each knowledge piece kk. [^6] Note that for those questions without knowledge information, we simply use an empty string as a keyword.
@@
[…]
Appendix F Data Process and Annotation
@@
F.1 Data Pre-process
The raw data range from HTML, photocopy, hand script, and plain text, and we conduct pre-processing to reduce the load of further annotation. We remove most HTML tags indicating irrelevant content of the question, such as alignment, color, etc. We reserve tags for underlines (<u> </u>), and we transfer several tagged styles, including bold, italic, and tabular data, into markdown format. For some complex tables that cannot be well converted, we save them as a screenshot picture after rendering with HTML.
For photocopy and hand script, we adopt OCR tools to convert text content, crop images, and figures, and integrate them into markdown. We further transcribe most of the math functions and chemistry structures into  format, with a small portion remaining as images.
F.2 Data Annotation
Figure 11: A screenshot of the main page of the data annotation platform.
We develop an online platform for the data annotation stage. The platform consists of text boxes for editing contents and regions for rendering the text to see the final appearance of the data as shown in Figure 11. We employ skilled human annotators and involve them as authors, primarily undergraduate students from top universities in China familiar with exam quizzes and markdown rules, to undertake this comprehensive task covering various aspects from formatting to semantic analysis:
@@ F.2 Data Annotation
  • Semantic Level. At this stage, annotators are advised to identify and correct both superficial errors (e.g., empty/duplicate choices, incomplete mathematical functions such as between \$32\$, \$3ˆ2\$, $\sqrt$$3\sqrt{2}$\$, $\frac${3}{2}\$) and more profound errors relating to factual accuracy and logical reasoning, such as content that is lacking or leads to inconclusive results. Those questions with profound errors will be dropped.
In Figure 12, we show several examples of complex formation and modification during the data annotation stage. The markdown, , and HTML format code remains for better formatting clarity.
@@
F.3 Data Post-process
To collect more challenging data for our benchmark, we adopt several data post-processing strategies:
@@ F.3 Data Post-process
  • Distillation. This is completed during our annotation process. We reduce assistance information so that the answer must depend on a more detailed analysis (refer to example (4) in Figure 13). In this way, we greatly enhance question difficulty.
  • Transformation. We randomly modify the questions, such as from single-choice to blank-filling (refer to example (2) in Figure 13), or convert certain kinds of single-choice questions into multiple-choice ones (refer to example (1) and (5) in Figure 13). Lots of single-choice questions have a list of options, and the choices are presented as a combination of those options where only one is correct. We transform those questions into multiple-choice questions where the options become new choices and the correct answer corresponds to the combinations. In this way, we successfully increase the scale of multiple-choice questions, improving the diversity of the questions.
@@
Figure 14: More examples of M ULTI .
Appendix G Case study
@@
Figure 15: Error analysis of an SI-SA question.
Figure 16: Error analysis of an MI-SA question.
Figure 17: Error analysis of an SI-MA question.
Figure 18: Error analysis of an SI-OP question.
Figure 19: Error analysis of an NI-MA question.